u-3(square root of u)-10=0 Hint: let x= square root of u, so x squared = u
Answer by edjones(8007)

u-3sqrt(u)-10=0
u-10=3sqrt(u)
u^2-20u+100=9u square each side.
u^2-29u+100=0
(u-25)(u-4)=0
u=25, u=4
u=25 is the only answer. If you check u=4 you will find that it is extraneous.
.
Ed